# Star Formation Histories of Dwarf Lenticular Galaxies

### v1

@@ 4.2.4 Background density
The SFHs of galaxies are known to be influenced by their environment (e.g., Kauffmann et al., 2004). In this study, we employed the background density () as a measure of galaxy environment. There are several methods to calculate background density (see Muldrew et al., 2012, for details), but for this analysis, we utilized the -th nearest neighbor method with . This method necessitates two constraints for selecting neighboring galaxies: the linking velocity () and the limiting magnitude (). We set the linking velocity to be 500 km s and the limiting magnitude as for local effect and for global effect. They are the r-band absolute magnitudes. The former corresponds to the luminosity to define a volume-limited sample for galaxies with redshifts less than and the latter is the defined in the luminosity function of the local universe (Ann et al., 2015).
@@
Figure 12 presents the cSFHs of dS0 galaxies grouped by their background density (). Similar to the approach used for physical parameters, we divided the galaxies into two groups based on the median and standard deviation () of the background density. The low density group consists of galaxies with , while the high density group includes galaxies with , where represents the median of the background density and is the standard deviation.
It is apparent that a distinction exists between the two groups in their cSFHs when grouped by background density, as constrained by . However, there is not a substantial difference observed for the two groups constrained by . This suggests that star formation in dS0 galaxies is likely more influenced by the local environment than by the global environment associated with large-scale structures. It is worth noting that the differences observed in the upper panel are statistically insignificant as the cSFHs fall within the boundary. The most pronounced difference in the cSFHs of the two density groups occurs around log() .
@@ 4.3 Quenching Epoch
Figure 13 displays the frequency distribution of quenching time () for 148 dS0 galaxies categorized by the presence of spiral arm features and nucleation. The term quenching time () was introduced by Weisz et al. (2014a, b, 2019) to quantify the cSFH of galaxies. It is defined as the lookback time at which 90% of the stellar mass has formed. The overall characteristics of the distribution resemble a Gaussian distribution with tails extending towards both the low and high ends. In particular, of dS0 galaxies exhibit greater than the lookback time of 10 Gyr. Similar distributions have also been observed in dSphs and dEs, with a higher fraction in dSphs (Seo & Ann, 2023).
As elucidated by Seo & Ann (2023), the extremely early quenching of early-type dwarf galaxies is likely due to their small mass, as low-mass galaxies are known to undergo quenching at earlier stages (Digby et al., 2019; Garrison-Kimmel et al., 2019; Joshi et al., 2021). It appears highly probable that dS0 galaxies with greater than the lookback time of 10 Gyr represent genuine primordial objects. The scarcity of such early-quenched dS0 galaxies suggests that the majority of dS0 galaxies may not be primordial in nature.
As anticipated due to the slight variations in the SFHs of dS0 galaxies with different morphological properties, there are discernible differences in values among dS0 galaxies with distinct morphologies. Concerning the presence of spiral arm features, no dS0 galaxies with spiral arms exhibit ) smaller than 9, while a small fraction of dS0 galaxies without spiral arms quenched at ) < 9. In contrast, there is no notable distinction in late quenching between dS0 galaxies with and without nucleation. It’s worth noting that the distribution of values appears somewhat narrower in dS0 galaxies with nucleation compared to those without nucleation, particularly when we disregard the high-end tail.
@@ 5 Discussion and Conclusions
In this study, we analyzed SDSS spectra to determine the star formation histories (SFHs) of 148 dS0 galaxies. To achieve this, we employed the population synthesis code STARLIGHT (Cid Fernandes et al., 2005). Our investigation has unveiled distinct characteristics within the SFHs of dS0 galaxies. At a lookback time of Gyr, these galaxies underwent an initial explosive starburst, followed by subsequent bursts of star formation peaking at a lookback time of 10 Gyr and Gyr. The peak occurring at a lookback time of Gyr primarily consists of older, metal-poor stars, while the peak at a lookback time of 2.5 Gyr predominantly comprises intermediate-metallicity stars. Notably, extremely metal-poor (Z=0.0001) stars are relatively rare during the early stages of galaxy formation, suggesting possible pre-enrichment influenced by Population III stars during the re-ionization period.
@@
Furthermore, our investigation suggests that stellar feedback plays a pivotal role in the early cessation of star formation in many dS0s, resulting in gaps in their SFHs. This early quenching effect is also observed in dSphs and dEs (Seo & Ann, 2023). Among the three types of early-type dwarfs, early quenching is most pronounced in dSphs, while dS0s exhibit an intermediate between dSphs and dEs.
@@
The star formation activity within dS0 galaxies can be categorized into three distinct periods. The first period exhibits two prominent peaks, one at a lookback time of Gyr and the other at a lookback time of 10 Gyr. The former marks the onset of an explosive starburst phase, contributing approximately 25% to the present-day stellar mass of dS0s. The latter epoch is notable for the generation of a majority of metal-poor stars.
@@
The second period spans from around the lookback time of 6.3 Gyr to Gyr, with the bulk of star formation occurring around the lookback time of Gyr. This phase predominantly gives rise to intermediate-metallicity stars. In dS0s, the stellar mass produced at the peak of the second period is comparable to that generated at the second peak of the first period. In the second period of star formation, the level of star formation activity is comparatively lower than that observed in the first period. However, stars with intermediate metallicity formed around the peak of the second period contribute of the current stellar mass. This substantial contribution is attributed to the stars formed during the peak of the second period of star formation. The third period of star formation recommences around the lookback times of Gyr after a phase of quiescent star formation. However, its contribution to the current stellar mass is deemed negligible.
@@
We have undertaken an investigation into the cSFHs of dS0 galaxies, taking into account various factors including morphology, stellar mass, u-r color, and background density. The cSFHs of dS0 galaxies reveal discernible dependencies on two crucial physical properties: stellar mass and u-r color. Notably, u-r color exerts a substantial influence on cSFHs. Conversely, the presence or absence of morphological features such as outer spiral arms and nucleation at the nuclei does not exhibit statistical significance. Additionally, while cSFHs in dS0 galaxies appear to evolve more rapidly in high-density areas, this correlation does not achieve statistical significance.
@@
The absence of significant dependence of cSFHs on morphological features, the outer spiral arms and nucleation, particularly during the early stages of galaxy formation, suggests that these morphological traits likely emerged after the initial period of star formation. This interpretation is supported by the observed differences in cSFHs that occur after the lookback time of Gyr. Furthermore, it is plausible that the outer spiral arms have minimal influence on star formation in the central regions of these galaxies, as these features are typically located outside the central star-forming regions.
@@
While the dependence of cSFHs on the background density () in dS0 galaxies is weaker compared to the influence of stellar mass, it is still a significant factor to consider. This is because some dS0 galaxies are likely the result of transformations from late-type galaxies, and morphology transformations are often driven by interactions with the environment. One indicator of this transformation is the presence of embedded structures in dS0 galaxies, reminiscent of late-type galaxies. A similar situation is observed in dEs, where a notable fraction of them also possess embedded disc features (Seo & Ann, 2022). It is not clear but we suggest that the presence of a second peak in the SFHs of dS0 galaxies and dEs is related to this morphology transformation. However, a substantial fraction of dS0s display early quenching, characterized by greater than a lookback time of 10 Gyr, suggesting their primordial origin. In contrast, dSph galaxies, which are mostly considered primordial objects, lack this second peak in their SFHs.
@@
The almost absence of extremely metal-poor (Z = 0.0001) stars in dS0 galaxies suggests the possibility of pre-enrichment of the gas from which these galaxies formed. This interpretation aligns with previous findings by Seo & Ann (2023), who proposed that the lack of extremely metal-poor stars in dEs and dSphs could be attributed to the rapid enrichment of metals by supernovae from Population III stars. This absence of extremely metal-poor stars is consistent with observations of metallicity in various astrophysical environments, including the Milky Way halo stars, local group dwarf spheroidal galaxies (Helmi et al., 2006), damped Ly absorption systems (Wolfe et al., 1998), and intervening intergalactic medium along sightlines to quasars (Cowie et al., 1998). The general absence of extremely metal-poor stars in the oldest stellar populations strongly suggests that the gases from which these stars formed were pre-enriched. This pre-enrichment, which resulted in the metallicity of the oldest stars reaching as high as Z = 0.0004, is consistent with numerical simulations (Tornatore et al., 2012; Wise et al., 2012) demonstrating the production of metals by Population III supernovae.
@@ Acknowledgements
H. B. Ann thanks Dr. Roberto Cid Fernandes for sending the source code of STARLIGHT.
@@
The original data underlying this article are available in the SDSS DR7. Additional data are available upon request.

### v2

@@ 4.2.3 Background density
The SFHs of galaxies are known to be influenced by their environment (e.g., Kauffmann et al., 2004). In this study, we employed the background density () as a measure of galaxy environment. There are several methods to calculate background density (see Muldrew et al., 2012, for details), but for this analysis, we utilized the -th nearest neighbor method with . This method necessitates two constraints for selecting neighboring galaxies: the linking velocity () and the limiting magnitude (). We set the linking velocity to be 500 km s-1 and the limiting magnitude as for local effect and for global effect. They are the r-band absolute magnitudes. The former corresponds to the luminosity to define a volume-limited sample for galaxies with redshifts less than and the latter is the defined in the luminosity function of the local universe (Ann et al., 2015). The choice of as 500 km s-1 is motivated by the peculiar velocity of the local universe (Peebles, 1979) as well as a detailed analysis of by Ann (2014).
@@
Figure 11 presents the cSFHs of dS0s grouped by their background density (). Similar to the approach used for physical parameters, we divided the galaxies into two groups based on the median and standard deviation () of the background density. The low density group consists of galaxies with , while the high density group includes galaxies with , where represents the median of the background density.
It is apparent that a distinction exists between the two groups in their cSFHs when grouped by background density, as constrained by . However, there is not a substantial difference observed for the two groups constrained by . This suggests that star formation in dS0s is likely more influenced by the local environment than by the global environment associated with large-scale structures. The most pronounced difference in the cSFHs of the two density groups occurs around the lookback time of 2 Gyr.
@@ 4.3 Quenching Epoch
Figure 12 displays the distribution of quenching time () for 148 dS0s categorized by the presence of spiral arms and nucleation. The quenching time (), introduced by Weisz et al. (2014a, b), quantifies the cSFH of a galaxy. It represents the lookback time at which 90% of the stellar mass has formed. The distribution of values is roughly Gaussian with tails at both ends. Notably, around 10% of dS0s exhibit exceeding 10 Gyr.
As anticipated due to slight variations in the SFHs of dS0s with different morphologies, there are discernible differences in values among dS0s with distinct morphologies. Concerning the presence of spiral arm features, no dS0 galaxy with spiral arms has a ) less than 9, while a small fraction of dS0s without spiral arms quenched at lookback time of Gyr. However, no significant difference in late-time quenching is observed between dS0s with and without nucleation.
In Figure 13, we present the distribution of for all dS0s, along with those of dSphs and dEs for comparison. As discussed in Seo & Ann (2023), the extremely early quenching observed in some early-type dwarf galaxies is likely due to their low mass. Galaxies with lower mass tend to quench at earlier stages (Digby et al., 2019; Garrison-Kimmel et al., 2019; Joshi et al., 2021). This aligns with the fact that dSphs have a larger fraction of early-quenched galaxies compared to dEs and dS0s. The distribution of dS0s bears a greater resemblance to that of dEs than to dSphs. Early-type dwarf galaxies with exceeding 10 Gyr represent genuine primordial objects.
@@ 5 Discussion and Conclusions
This analysis of the star formation histories (SFHs) of 148 dS0s, using SDSS spectra and STARLIGHT (Cid Fernandes et al., 2005), revealed distinct characteristics. A prominent initial burst of star formation occurred at a lookback time of around 14 Gyr, followed by subsequent bursts that peaked at around 10 Gyr and 2.5 Gyr. The 10 Gyr peak was dominated by metal-poor stars, while the 2.5 Gyr peak was rich in stars with intermediate metallicity. Interestingly, extremely metal-poor stars (Z=0.0001) are scarce during early galaxy formation, suggesting potential pre-enrichment by Population III stars during the re-ionization era.
@@
Our investigation further suggests that stellar feedback plays a crucial role in the early cessation of star formation in many dS0s, resulting in gaps in their SFHs. This early quenching effect is also observed in dSphs and dEs (Seo & Ann, 2023). However, among the three types of early-type dwarf galaxies, early quenching is most pronounced in dSphs, while dS0s exhibit an intermediate level compared to dSphs and dEs.
@@
The star formation activity within dS0s can be divided into three distinct periods. The first period exhibits two prominent peaks, one at the lookback time of 14 Gyr and the other at the lookback time of 10 Gyr. The first peak marks the onset of a major starburst phase, responsible for 27% of the present-day stellar mass in dS0 galaxies. The second peak represents another significant epoch for the formation of metal-poor stars.
@@
The second period spans from between the lookback time of 6 Gyr to 1 Gyr, with the bulk of star formation occurring at the lookback time of 2.5 Gyr. This period predominantly gives rise to intermediate-metallicity stars. In dS0s, the stellar mass produced at the peak of the second period is comparable to that generated during the earlier peak at 10 Gyr. In the second period of star formation, the level of star formation activity is comparatively lower than that observed in the first period. However, stars with intermediate metallicity formed around the peak of the second period contribute 30% of the current stellar mass. The third period of star formation recommences around the lookback times of Gyr after a phase of quiescent star formation. However, its contribution to the current stellar mass is negligible.
@@
We have undertaken an investigation into the cSFHs of dS0s, taking into account various factors including morphology, stellar mass, and background density. The cSFHs of dS0s reveal discernible dependencies on physical properties: stellar mass and local background density. In constrast, the presence or absence of morphological features such as outer spiral arms and nucleation do not exhibit statistical significance. Additionally, it is also apparent that dS0s appear to form stars more rapidly in high-density environments.
@@
The lack of significant dependence of cSFHs on morphological features, the outer spiral arms and nucleation, particularly during the early stages of galaxy formation, suggests that these morphological traits likely developed after the initial period of star formation. This is supported by the observed differences in cSFHs that occur after the lookback time of Gyr. Additionally, the outer spiral arms have minimal influence on star formation in the central regions of these galaxies, as these features are typically located outside the central star-forming regions.
@@
While the dependence of cSFHs on the background density () is weaker than stellar mass in dS0s, it remains a significant factor. This is because some dS0s might have transformed from late-type galaxies, often driven by environmental interactions. One indicator of this transformation is the presence of embedded structures in dS0 galaxies, reminiscent of late-type galaxies. A similar situation is observed in dEs, where a notable fraction of them also possess embedded disc features (Seo & Ann, 2022). We propose that the second peak in the SFHs of dS0s and dEs might be linked to this potential transformation. However, a significant fraction of dS0s exhibit early quenching ( > 10 Gyr), suggesting a primordial origin. In contrast, dSphs, mostly considered primordial objects, lack this second peak.
@@
The scarcity of extremely metal-poor (Z = 0.0001) stars in dS0s hints at pre-enrichment of the gas that formed these galaxies. This aligns with findings in dEs and dSphs (Seo & Ann, 2023), where the absence of such stars is attributed to the rapid metal enrichment by Population III supernovae. This lack of extremely metal-poor stars is consistent with observations across various environments, including the Milky Way halo stars, Local Group dwarf spheroidals (Helmi et al., 2006), damped Ly absorption systems (Wolfe & Prochaska, 1998), and the intergalactic medium towards quasars (Cowie & Songaila, 1998).Simulations (Wise et al., 2012) also support pre-enrichment, potentially leading to metallicities as high as Z = 0.0004 for the oldest stars.
@@ Acknowledgements
HBA thanks the anonymous reviewer for their valuable comments, which significantly improved the paper. He also thanks Dr. Roberto Cid Fernandes for providing STARLIGHT.
@@
We provide the basic output of STARLIGHT as supplemented materials and additional data are available upon request.
Appendix A Cumulative Star Formation Histories of 148 dS0 galaxies
Figure A1 shows the cSFHs of the 148 dS0s. A small fraction () of dS0s exhibit very early quenching, while a similar fraction of dS0s show very delayed star formation. The SFHs of the 148 dS0s are provided in tabular form as supplementary data.